Could someone please help me with these?
Evaluate each of the following give:
f(x)= -7x,g(x) = 4x-2 and h (x)= 2x(squared).
a) (f+g)(x)
b) (h-f)(x)
c) (h+f)(6)
d) (f o h)(x)

Thank you

Answer by Alan3354(69443)

Could someone please help me with these?
Evaluate each of the following give:
f(x)= -7x,g(x) = 4x-2 and h (x)= 2x(squared).
a) (f+g)(x)
Add f(x) and g(x)
--> -7x + 4x-2
= -3x - 2
====================
b) (h-f)(x)
Similar to a.
Subtract f(x) from h(x)
=====================
c) (h+f)(6)
Find h(6) and f(6)
h(x)= 2x^2
h(6) = 2*6^2 = 72
---
Find f(6), then add them.
===========================
d) (f o h)(x)
h(x) = 2x^2
--
f(x)= -7x
Sub 2x^2 for x.